# ConfuGuard: A Low False-Positive System for Detecting Malicious and Stealthy Typosquatting Threats in Package Registries

### v2

@@ 4.1.2 Extending the Taxonomy to New SPRs
Prior work has examined malicious naming strategies in only a few SPRs. Our industry partner’s analysts have studied package confusion attacks in many other SPRs, including all those in Table 1. We discussed Neupane et al.’s taxonomy with them, and compared it to the real-world attack patterns they have observed. Through this dialogue, we identified four additional mechanisms for package confusion attacks. Definitions follow; see section 2.2 for examples.
@@
  • Domain Confusion. In SPRs where package names include URLs (golang), attackers may register domains resembling those of trusted mirrors or proxies.
  • Command Squatting. Registering package names that mimic command-line options in other packages or utilities.
@@ 4.2 Metadata Features of Package Confusion
As discussed in section 2.3, almost all package confusion detectors make use only of the names of packages. While this allows those systems to avoid costly content analysis, it leads to high false positive rates. We propose to integrate metadata to distinguish true from false positives. Although metadata signals can be circumvented, our approach aims to raise the attacker’s cost, reflecting the typical cat-and-mouse dynamics in open package ecosystems [28].
Table 3: Overview of the 13 metadata-based verification rules. Each rule includes a description of its purpose and the specific implementation steps taken to verify flagged packages. The final three rules (R12–R14) were added as part of our refinement process based on further observed false-positive patterns after we deployed our system in production.
Rule Description Implementation
@@
R1: Obviously unconfusing Determine if a package name conveys an intentional, brand-specific identity (e.g., catplotlib), clearly differentiating it from deceptive imitations. Cross-reference flagged names with known legitimate projects and use LLM analysis to assess whether the naming convention is deliberate and unambiguous for developers.
R2: Distinct Purpose Distinguish packages with different functionalities, even if names are superficially similar (e.g., lodash-utils vs. lodash). Extract package descriptions and calculate semantic similarity using TF-IDF cosine scores. A score below 0.5 indicates distinct purposes, reducing suspicion of deception.
@@
R3: Fork Package Detect benign forks sharing near-identical code or metadata with a popular package. Compare README files, version histories, and file structures for high overlap. Similarities without malicious edits suggest harmless forks.
@@
R4: Active Development/Maintained Determine if packages are frequently updated or actively maintained by multiple contributors, which are less likely to have malicious intent. Retrieve metadata for the last update, commit history, and version count. Classify packages with recent updates (e.g., within 30 days) or more than five versions as legitimate.
@@
R5: Comprehensive Metadata Identify packages missing critical metadata elements, such as licenses, maintainers, or homepages, which are typical of legitimate projects. Check for the presence of licenses, contact details, and repository links.
@@
R6: Overlapped Maintainers Distinguish legitimate extensions or rebrands by verifying if the flagged package shares maintainers with the legitimate one. Match maintainer identifiers (e.g., email, GitHub handle) between flagged and legitimate packages. Overlapping maintainers suggest legitimate intent.
@@
R7: Adversarial Package Name Filter out name pairs with significant length differences, as these often indicate unrelated projects rather than covert mimicry. Compare string lengths of flagged and legitimate package names. A difference exceeding 30% indicates likely unrelated naming.
@@
R8: Well-known Maintainers Trust packages maintained by reputable and recognized authors/organizations. Leverage knowledge in LLM training data to identify if a maintainer is trustworthy in the community.
@@
R9: Clear Description Verify that each package includes a README or equivalent documentation detailing its purpose, usage, and key features. Analyze repository metadata to ensure there is a clear, summarized description of the package; packages lacking such documentation should be flagged for further review.
@@
R10: Has Malicious Intent Identify packages that deliberately mimic legitimate ones through near-identical descriptions, or have obviously suspicious content. Use LLM to analyze the package name and description to triage whether there is obvious malicious intent in the package.
@@
R11: Experiment/Test Package Identify packages that are used for test or experiment purposes only. Use LLM to analyze package descriptions and determine whether it is an experiment/test.
@@
R12: Package Relocation Account for legitimate package relocations, common in hierarchical registries like Maven. Parse metadata files (pom.xml) for <relocation> tags or analogous fields. Identify and ignore renamed or migrated projects.
@@
R13: Organization Allowed List Prevent false positives by excluding packages published by trusted or verified organizations. Maintain an allowedlist of approved organizations. If a flagged package is published under an allowed organization (e.g., @oxc-parser/binding-darwin-arm64), it should be considered legitimate, comparing to binding-darwin-arm64.
@@
R14: Domain Proxy/Mirror Account for legitimate proxies or mirrors in the ecosystem, common in hierarchical registries like Golang. Maintain an allowed list of recognized domains that serve as proxies or mirrors. If a given package is published under a valid domain (e.g., gopkg.in/go-git/go-git), consider it legitimate when compared to its primary source (github.com/go-git/go-git).
@@ 4.2.1 Method
We randomly sampled 626 packages from the original 640,482 false positives reported by Neupane et al. [39], which have suspicious package names but are not flagged as malware in their respective security analyses. This sample size gives a confidence level of 99% with a margin of error of 5% on the resulting distribution of causes.
To identify possible metadata signals in the existing package confusion data, we began by having two researchers — experts in software supply chains — analyze 200 of these packages. They analyzed each package’s metadata (READMEs, maintainers, versions, etc.). Each analyst independently proposed possible features based on this analysis (codebook) and then they discussed this codebook together to reach agreement on terms and definitions. To test valididty, they then independently applied this codebook to the 200 packages and measured agreement using Cohen’s Kappa [18]. The initial inter-rater reliability was 0.6 (“substantial” [35]). The researchers subsequently discussed to resolve discrepancies and refine their analysis. Through discussion, they reached consensus on measurable attributes that could indicate malicious intent or the possibility of a stealthy attack.
@@ 4.2.2 Results
Of the 626 packages sampled, 601 were still accessible. Among these, we identified 464 benign packages (77.2%) and what we believe are 137 stealthy attacks (22.8%). The progression of our results and the manually labeled dataset are available in our artifact (Data Availability and Research Ethics). We identified 11 metadata features of the packages used in confusion attacks, as well as benign signals. These attributes include factors such as a distinct purpose, adversarial package naming, and the comprehensiveness of available metadata. See Table 3 for the features leveraged in ConfuGuard.
Figure 3: Overview of the ConfuGuard design with five primary components. The main novelty of ConfuGuard is the techniques of Part 4 and 5. We also improve on prior work in Part 2. and Part 1 (DB) can be constructed using public registry APIs. Part 3’s approach for popular packages is from prior work, adapted for the broader set of SPRs we handle [56, 39]. The red texts indicate the evaluation questions (EQs).
@@ 5 ConfuGuard Design and Implementation
We introduce ConfuGuard, our novel detector for package confusion attacks. ConfuGuard is designed to meet the system requirements (section 3.3). Addressing the insights from our empirical study (section 4), ConfuGuard integrates both syntactic and semantic name analysis, hierarchical naming checks, and metadata-based verification to enhance threat detection and mitigation. ConfuGuard is intended for backend use, in package registries or similar platforms, and so it prioritizes accuracy (Req1) over latency (Req2).
Figure 3 shows the five main components of ConfuGuard:
  1. 1.
    A metadata database, ensuring real-time awareness of new and evolving packages (Req4).
  2. 2.
    A fine-tuned embedding database to capture domain-specific semantic name similarities (Req1,3).
  3. 3.
    Using popularity metrics to protect high-value targets (Req2,3).
  4. 4.
    Checking for syntactic and semantic confusion strategies (Req2).
  5. 5.
    Filtering out benign packages using package metadata (Req1,4).
The first three components provide infrastructure to support our analysis (components 4 and 5). Next, we detail the rationale and implementation of each part of ConfuGuard.
@@ 5.1.1 Rationale
SPRs grow rapidly, and package confusion attacks are initiated regularly. Comprehensive and regularly updated metadata ingestion enable early threat detection.
@@ 5.1.2 Approach
ConfuGuard relies on a database with package names, version histories, commit logs, license info, maintainer records, and other publicly-accessible metadata, all updated on a weekly basis. Regular updates reduce concerns about stale data and ensure our results are up to date. We specifically use our industry partner’s private database, which consolidates metadata from NPM, PyPI, RubyGems, Maven, Golang, and Hugging Face.
@@ 5.2.2 Approach
Like prior works [56, 39], we operationalize trust in terms of popularity. Software engineers commonly make use of popularity signals as a proxy for trustworthiness [15], and thus attackers choose names similar to popular packages. Our specific popularity measure depends on the SPR. Many SPRs — in our case, NPM, PyPI, RubyGems, and Hugging Face – offer weekly or monthly download counts. Our other two supported SPRs, Maven and Golang, do not publish download metrics. For these, we make use of the ecosyste.ms database, which estimates popularity using indicators like stargazers, forks, and dependent repositories [38]. The thresholds were set based on production experience. We chose 5,000 weekly downloads for NPM, PyPI, and RubyGems; 1,000 for Hugging Face due to its costly usage; and an avg_ranking score of 10 for Maven and Golang. We then adjusted the Golang thresholds to 4 to balance production latency.
Adversaries may inflate the popularity statistics of their packages [24], which could cause a malicious (or decoy) package to become trusted. If our partner’s analyst team flags any package as suspicious (section 5.6), it is removed from the list of trusted packages.
As a secondary definition of trust, we analyzed the command-line interfaces of tooling for each SPR, e.g., the “npm” and “mvn” commands. These interfaces contain built-in keywords and commands which a careless engineer might confuse or mis-type for package names, e.g., mistaking npm i help (install a package named help) for npm i --- -help (help command). All such keywords are included as trusted resources.
@@ 5.3.1 Rationale
Detecting maliciously similar names requires accurately capturing subtle lexical, syntactic, and semantic variations. Traditional Levenshtein edit distance methods often fail to account for domain-specific semantic nuances (e.g., meta-llama vs. facebook-llama), while generic embedding models can introduce inaccuracies, resulting in higher false-positive or false-negative rates. Prior work apply embedding for word-wise semantic similarity check [39], while we would also like to use embedding for the whole name to support efficient neighbor search (section 5.4). Robust, fine-tuned embeddings address these shortcomings by providing enhanced semantic sensitivity, reducing erroneous alerts. Moreover, this embedding approach is also generic and unified to support various naming convention of SPRs.
@@ 5.3.2 Approach
An embedding fine-tuned on real package names enhances semantic sensitivity, enabling the detection of adversarial or suspicious names. This capability is integral to assessing the risk associated with a package and serves as a cornerstone of our intent-centric approach. Following [39], we build upon FastText [14], starting with the pre-trained model pre-trained on cc.en.300.bin, and fine-tune it using all (totally $\sim$9.1 million) package names extracted from the metadata database in Nov. 2024 (§5.1). By fine-tuning, our version is expected to better capture domain-specific subwords.
We then use the fine-tuned embedding model to create an embedding vector database utilizing the vector format provided by pgvector due to its efficient vector operations for databases [9]. Given a package name, we remove its delimiters, pass the concatenated string to the embedding model, and store the extracted embedding in our database. We note that for hierarchical names from Golang (domain/author/repository), Maven (groupId:artifactId), and Hugging Face (author/model), we split names to create both author and package identifier embeddings per package. This approach mitigates high false-positive rates that occur when a single embedding causes similar packages from the same organization or author to be incorrectly classified as confusing in section 5.4.
The complete embedding database occupies 24 GB, with each embedding vector corresponding to a single package name (or its author name and package identifier). This setup not only facilitates rapid query-based lookups but also supports subsequent steps in package neighbor searching. Visualization of the extracted embedding vectors is available in our artifact (Data Availability and Research Ethics).
@@
Having obtained the appropriate infrastructure (Parts 1–3), we must now detect package confusion threats efficiently yet accurately. In Part 4, we efficiently flag possible threats, and then in Part 5 we filter out likely-benign ones.
@@ 5.4.2 Approach
Trusted Resources Selection
We compare untrusted resources (i.e., unpopular packages) against all names in the popular list. For packages in the trusted resources list (i.e., popular packages), we compare it only with “more” trusted packages — for popularity, we use download rates $\geq$10x higher or ecosyste.ms scores $\geq$2x higher.
Neighbor Search (Package Name Similarity Search)
We perform a distance search between each resource and the set of trusted resources. Syntactic similarity is measured using Levenshtein distance with a threshold of 2, while semantic similarity is determined using embedding distance. For semantic similarity, we apply a cosine similarity threshold of 0.93, as suggested in section 6.2. For hierarchical names, we adjust the similarity thresholds to 0.99 for package identifiers and 0.9 for author names to better detect compound squatting attacks. To efficiently measure semantic similarity, we leverage the Approximate Nearest Neighbors method with the HNSW index [34] in PostgreSQL. We opted for HNSW over IVFFlat based on benchmarking results that show faster search speeds and minimal vector perturbation [58]. By partitioning the embedding space into multiple clusters, the HNSW index limits distance computations to a smaller subset of candidate packages, rather than exhaustively comparing all pairs (Req3).
Similarity Sort and Most Similar Neighbor(s)
Inspired by Neupane et al. [39], we define a similarity function that ranks neighbors based on Levenshtein distance, n-gram similarity, phonetics, substring matching, and fuzzy ratio. This sorting ensures the system accurately identifies the most likely attack target. If the previous steps produce any nearby neighbors, we consider these as possible attack targets. We treat the nearest neighbors as the most likely targets. The suspicious package and these nearest neighbors are propagated to the benignity check (section 5.5). Currently we consider only the two nearest neighbors, balancing accuracy against speed.

### v3

@@ 4.1.2 Extending the Taxonomy to New SPRs
Prior work characterized malicious naming strategies in only a small set of SPRs, leaving gaps for ecosystems with hierarchical namespaces and platforms for hosting pre-trained AI models. In collaboration with our industry partner’s security analysts, who routinely triage package confusion incidents across all SPRs in Table˜1, we reviewed Neupane et al.’s taxonomy against their field observations. Through this dialogue, we identified three additional mechanisms for package confusion attacks. Definitions follow; see Table˜2 for examples.
@@
  • Domain Confusion. In SPRs where package names include URLs (e.g., golang), attackers may register domains resembling those of trusted mirrors or proxies.
@@ 4.2 Metadata Features of Package Confusion
As discussed in section˜2.3, almost all package confusion detectors make use only of the names of packages. While this allows those systems to avoid costly content analysis, it leads to high false positive rates. We propose to integrate metadata to distinguish true from false positives. Although metadata signals can be circumvented, our approach aims to raise the attacker’s cost, reflecting the typical cat-and-mouse dynamics in open package ecosystems [37].
Table 3: Overview of the 15 metadata-based verification rules. Each rule lists its goal, the concrete implementation & metadata used (with inspected metadata fields italicized), and the mapped LLM directive. In the LLM Directive column, “—” denotes a rule implemented purely with metadata checks and no LLM prompt. The final four rules (R12–R15) were added during deployment to reduce false positives.
Rule Description Implementation & metadata used LLM Directive
@@
R1: Obviously Unconfusing Determine if a package name conveys an intentional, brand-specific identity (e.g., catplotlib), clearly differentiating it from deceptive imitations. Cross-reference flagged package names with known legitimate projects and use LLM analysis to assess whether the naming convention is deliberate and unambiguous for developers. obvious_not_typosquat
R2: Distinct Purpose Distinguish packages with different functionalities, even if names are superficially similar (e.g., lodash-utils vs. lodash). Pass both package descriptions to the LLM; if the model reasons they serve genuinely different use cases or unique functionality, mark as distinct purpose, otherwise flag as possible impersonation. has_distinct_purpose
@@
R3: Fork Package Detect benign forks sharing near-identical code or metadata with a popular package. Compare README files, version histories, and file structures for high overlap. Similarities without malicious edits suggest harmless forks. is_fork
@@
R4: Active Development/Maintained Determine if packages are frequently updated or actively maintained by multiple contributors, which are less likely to have malicious intent. Retrieve last-update timestamp, commit history, and version count. Classify packages with recent updates (e.g., within 30 days) or more than five versions as legitimate.
@@
R5: Comprehensive Metadata Packages missing critical metadata elements are suspicious. Check for licenses, maintainer information, and repository links. no_readme
@@
R6: Overlapped Maintainers Distinguish legitimate extensions or rebrands by verifying if the flagged package shares maintainers with the legitimate one. Match maintainer identifiers (e.g., email) between flagged and legitimate packages. Overlapping maintainers suggest legitimate intent. is_known_maintainer
@@
R7: Adversarial Package Name Filter out name pairs with significant length differences, as these often indicate unrelated projects rather than covert mimicry. Compare flagged and legitimate package name lengths. A difference exceeding 30% indicates likely unrelated naming. is_adversarial_name
@@
R8: Well-known Maintainers Trust packages maintained by reputable and recognized authors/organizations. Leverage knowledge in LLM training data to identify reputable maintainers in the community. is_known_maintainer
@@
R9: Clear Description Legitimate packages document purpose and usage. Analyze repository metadata to ensure there are clear package description and README; packages lacking such documentation should be flagged for further review. no_readme
@@
R10: Has Malicious Intent Identify packages that deliberately mimic legitimate ones through near-identical descriptions, or have obviously suspicious content. Use LLM to analyze the package name and package description to triage whether there is obvious malicious intent in the package. has_suspicious_intent
@@
R11: Experiment/Test Package Detect packages that are used for test or experiment purposes only. Use LLM to analyze package description and determine whether the package is for experimental use. is_test
@@
R12: Package Relocation Account for legitimate package relocations, common in hierarchical registries like Maven and Golang. Parse project configuration files (e.g., look for <relocation> tag inside pom.xml). Identify and ignore renamed or migrated projects. is_relocated_package
@@
R13: Organization Allowed List Prevent false positives by excluding packages published by trusted or verified organizations. Maintain a approved organizations list. If a flagged package is published under an allowed organization (e.g., @oxc-parser/binding-darwin-arm64), it should be considered legitimate, comparing to binding-darwin-arm64.
@@
R14: Domain Proxy/Mirror Account for legitimate proxies or mirrors in the ecosystem, common in hierarchical registries like Golang. Maintain a recognized domains list that serve as proxies or mirrors. If a given package is published under a valid domain (e.g., gopkg.in/go-git/go-git), consider it legitimate when compared to its primary source (github.com/go-git/go-git).
R15: Registered Prefix Leverage NuGet’s prefix reservation: legitimate publishers can reserve name prefixes (e.g., Newtonsoft.Json.*). Nuget offers a verified field to indicate if a package name has a registered prefix. A verified package is considered legitimate.
@@ 4.2.1 Method
We randomly sampled 626 packages from the original 640,482 false positives reported by Neupane et al. [49], which have suspicious package names but are not flagged as malware in their respective security analyses. This sample size gives a confidence level of 99% with a margin of error of 5% on the distribution.
To identify possible metadata signals in the existing package confusion data, we began by having two researchers analyze 200 of these packages. They analyzed each package’s metadata (e.g., READMEs, maintainers, and version timelines). Each researcher independently proposed possible features based on this analysis (codebook) and then they discussed this codebook together to reach agreement on terms and definitions. To test validity, they then independently applied this codebook to the 200 packages and measured agreement using Cohen’s Kappa [15]. The initial inter-rater reliability was 0.6 (“substantial” [45]). The researchers then convened to resolve discrepancies and refine their analysis. Through discussion, they reached a consensus on measurable attributes that could indicate malicious intent or the possibility of a stealthy attack.
@@ 4.2.2 Results
Of the 626 packages sampled, 601 were still accessible. Among these, we identified 480 benign packages (79.9%) and what we believe are 121 stealthy attacks (20.1%). The progression of our results and the manually labeled dataset are available in our artifact (Data Availability). From this analysis, we distilled 11 package metadata-based rules that differentiate genuine confusion attacks from benign look-alikes. These attributes include factors such as a distinct purpose, adversarial package naming, and the comprehensiveness of available metadata. See Table˜3 for ConfuGuard’s rules.
@@ 5 ConfuGuard Design and Implementation
We introduce ConfuGuard, our detector for package confusion threats and attacks. ConfuGuard is designed to meet the system requirements defined in section˜3.3. Leveraging the insights from our empirical study (section˜4), ConfuGuard integrates syntactic and semantic package name analysis, hierarchical naming checks, and verifications based on package metadata to enhance threat detection.
Figure˜3 shows the components and parts of ConfuGuard:
  1. Component I.
    Infrastructure:
[…]
      A package metadata database, ensuring real-time awareness of new and evolving packages (Req4).
[…]
      Popularity metrics to protect high-value targets (Req2,3).
[…]
      A fine-tuned embedding model and database to capture domain-specific semantic name similarities (Req1,3).
  2. Component II.
    System Analysis:
[…]
      Syntactic and semantic confusion checks (Req2).
[…]
      Filtering out benign packages using metadata (Req1,4).
  3. Component III.
    Human Analysis:
[…]
      Forwarding possible threats to security analysts (Req1,4).
[…]
@@ 5.1.1 Rationale
Package confusion attacks are initiated regularly. Up-to-date metadata facilitates threat detection.
@@ 5.1.2 Approach
ConfuGuard relies on a database with package names, version histories, commit logs, license info, maintainer records, and other publicly accessible metadata, all updated on a weekly basis. Regular updates reduce concerns about stale data and ensure our results are up to date. We specifically use our industry partner’s private database, which consolidates metadata from NPM, PyPI, RubyGems, Maven, Golang, Hugging Face, and NuGet.
@@ 5.2.2 Approach
Like prior works [68, 49], we operationalize trust in terms of popularity. Software engineers commonly make use of popularity signals as a proxy for trustworthiness [10], and thus attackers choose names similar to popular packages. Our specific popularity measure depends on the SPR. Many SPRs — in our case, NPM, PyPI, RubyGems, Hugging Face, and NuGet – offer weekly or monthly download counts. Our other two supported SPRs, Maven and Golang, do not publish download metrics. For these, we make use of the ecosyste.ms database, which estimates popularity using indicators like stargazers, forks, and dependent repositories [48]. The thresholds were set based on production experience and costs. We initially chose 5,000 weekly downloads for NPM, PyPI, RubyGems, and NuGet; 1,000 for Hugging Face (normalized for its lower download rate [33, 32]); and an avg_ranking score of 10 for Maven and Golang. We then adjusted the Golang thresholds to 4 to balance production latency.
Adversaries may inflate the popularity statistics of their packages [30], which could cause a malicious (or decoy) package to become trusted. If our partner’s security analysts flag any package as suspicious (section˜5.6), it is removed from the list of trusted packages.
@@ 5.3.1 Rationale
Detecting maliciously similar names requires knowledge of lexical, syntactic, and semantic variations. Simple Levenshtein edit distance methods omit domain-specific semantic nuances (e.g., meta-llama vs. facebook-llama on Hugging Face), while generic embedding models can introduce inaccuracies that result in higher false positive or false negative rates. Robust, fine-tuned embeddings address these shortcomings by providing enhanced semantic sensitivity, reducing erroneous alerts. Prior work applies embedding for word-wise semantic similarity checks [49]. We also use embedding for the whole name to support efficient neighbor search (section˜5.4). A generic embedding approach also supports various package naming conventions, e.g., differences by SPR.
@@ 5.3.2 Approach
An embedding model fine-tuned on real package names enhances semantic sensitivity, enabling the detection of adversarial or suspicious names. This capability is integral to assessing the risk associated with a package and serves as a cornerstone of our intent-centric approach. Following [49], we build upon FastText [9], starting with the pre-trained model pre-trained on cc.en.300.bin, and fine-tune it using all $\sim$9.1 million package names in our package metadata database as of Nov. 2024 (§5.1). By fine-tuning, we expect to better capture domain-specific semantics.
We then use the fine-tuned embedding model to create an embedding database. Given a package name, we remove its delimiters and compute the embedding of the concatenated string. For hierarchical names from Golang (domain/author/repository), Maven (groupId:artifactId), and Hugging Face (author/model), we split names to create distinct author and package identifier embeddings. This approach reduces false positives wherein similar packages from the same organization or author are incorrectly classified as confusing (section˜5.4).
The resulting embedding database occupies 24 GB. We use the vector format provided by pgvector due to its efficient vector operations for databases [53]. This setup facilitates rapid query-based lookups and supports subsequent steps in package neighbor searching. Our artifact has visualizations of the embedding vectors (Data Availability).
@@
Having obtained the appropriate infrastructure, we must now detect package confusion threats efficiently yet accurately. In this part, we focus on efficiently flagging possible package confusion threats.
@@ 5.4.2 Approach
To support the threshold choices and parameter settings used in ConfuGuard for detecting package confusion attacks, we studied the naming characteristics in these attacks. We analyzed the 1,239 legitimate and typosquatted package name pairs from NPM, PyPI, and RubyGems provided by Neupane et al. [49]. Our analysis examined two key aspects of legitimate and typosquatted package name similarities: lexical distance using Levenshtein distance and semantic similarity using cosine similarity. For semantic similarity measurements, we used cosine similarities of package names computed from our fine-tuned FastText embeddings (section˜5.3).
We confirmed that real-world package confusion attacks involve small syntactic changes and high semantic similarity. The median Levenshtein distance between typosquatted and legitimate package names is 1, while the mean distance is 2.54. Additionally, 77.7% of all package pairs exhibit a distance of 2 or less. Among the analyzed package pairs, the mean cosine similarity was 0.95, and the median attained the maximum possible value of 1.
Trusted Resources Check We compare untrusted resources (i.e., unpopular packages) against all packages on the popular list. For trusted resources (i.e., popular packages), we restrict comparisons to “more” trusted packages. A package is considered more trusted than a given trusted package if it has a download rate at least 10 times higher or an ecosyste.ms score at least twice as high — following the criteria previously used by our industry partner.
Neighbor Search (Package Name Similarity Search) We perform a distance search between each resource and the set of trusted resources. Syntactic similarity is measured using Levenshtein distance (threshold=2=2= 2), seeded by our empirical study and refined during deployment. For semantic similarity, we apply a cosine similarity threshold of 0.93, likewise initialized from the empirical baseline and finalized through deployment feedback. As a special case for hierarchical names, we tighten the similarity thresholds to 0.99 for package names and 0.9 for author names to better detect “compound squatting” attacks (cf. Table˜6); these tighter bounds reflect security analysts’ observations during our field tests.
To efficiently measure semantic similarity, we leverage the Approximate Nearest Neighbors method with the HNSW index [44] in PostgreSQL. We opted for HNSW over IVFFlat based on benchmarking results that show faster search speeds and minimal vector perturbation [70]. By partitioning the embedding space into multiple clusters, the HNSW index limits distance computations to a smaller subset of candidate packages, rather than exhaustively comparing all pairs (Req2).
Similarity Sort and Most Similar Neighbor(s) Inspired by Neupane et al. [49], we define a similarity function that ranks neighbors based on Levenshtein distance, n-gram similarity, phonetics, substring matching, and fuzzy ratio. We treat the nearest neighbors as the most likely targets. The suspicious package and these nearest neighbors are propagated to the benignity check (section˜5.5).
In production, similarity sorting often returns multiple packages with near-perfect scores. section˜6.6 shows that adding more neighbors substantially increases latency and the cost of the benignity check, while delivering minimal accuracy gains. Based on our industry partner’s feedback, we restrict retrieval to the top-2 neighbors to optimize the trade-off between accuracy and speed.